OpenStudy (anonymous):

2x+6y=24 and 5x-2y=9 soving by subsitution

OpenStudy (marigirl):

there are numerous ways to start, so lets just begin somewhere! lol first let us take the first equation and x the subject. 2x+6y=24 2x =24-6y... so what would be the next line? x =

OpenStudy (anonymous):

x=24-6y/2

OpenStudy (marigirl):

yes lets simplify to y=12-3y... .happy with this?

OpenStudy (marigirl):

now lets substitute x=12-3y in to the second equation 5x-2y=9. basically we are replacing the x variable with 12-3y 5(12-3y)-2y=9

OpenStudy (marigirl):

5(12-3y)-2y=9 lets expand 60-15y-2y=9 .. can u solve for y now and tell me what value u get?

OpenStudy (anonymous):

17 y=51 Can you find y now?

OpenStudy (anonymous):

\[ \frac {51}{17}=3 \]